[38-(4+6)÷2] - [1+24÷3]
1
Expert's answer
2011-07-20T08:59:05-0400
[38-(4+6)÷2] - [1+24÷3] = [38 - 10÷2] - [1 + 8] = [38& - 5] - 9 = 33 - 9 = 24